# Physics-based Full-band GaN HEMT Simulation Suggests Upper Bound of LO Phonon Lifetime

### v1

@@ III.3 Electron and Phonon Temperatures
The temperature variations in Fig. 10 also reveal certain characteristics of acoustic phonon lattice heating. In contrast to the electron temperature profile, the lattice temperature exhibits a more gradual spatial variation, increasing smoothly along the channel due to the slower energy transfer from electrons to acoustic phonons mediated by LO phonon decay. The lattice temperature’s dependence on $\tau_{\mathrm{LO}}$ also appears to be in contrast to the electron temperature. Under nonequilibrium steady-state conditions, the power supplies inject energy into the device as $V_{\mathrm{DS}}I_{\mathrm{D}}$, which first heats the mobile electrons through Joule heating. The hot electrons, in turn, transfer their energy to the LO phonons through scattering, which subsequently decay into acoustic phonons that ultimately diffuse heat into the heat sinks. Increasing the LO phonon lifetime, $\tau_{\mathrm{LO}}$, strengthens the hot-phonon bottleneck where the long-lived LO phonons accumulate in the channel, forming a nonequilibrium population that repeatedly reabsorbs energy from the hot electrons. This elevates $T_{\mathrm{LO}}$, suppresses electron mobility, and reduces the overall current and power dissipation $V_{\mathrm{DS}}I_{\mathrm{D}}$. The resulting drop in $V_{\mathrm{DS}}I_{\mathrm{D}}$ leads to a lower rate of energy that needs to be carried away by the acoustic phonons, thereby reducing the maximum $T_{\mathrm{A}}$. This establishes a negative feedback mechanism, where we observe that increased $\tau_{\mathrm{LO}}$ leads to stronger electron heating but weaker lattice self-heating across the GaN channel. This explains the opposing trends of the peak lattice temperature and LO phonon temperature observed in Fig. 11, where increasing $\tau_{\mathrm{LO}}$ causes the peak $T_{\mathrm{LO}}$ to rise while the peak $T_{\mathrm{A}}$ decreases. Therefore, reduced lattice self-heating with longer phonon lifetime originates not directly from phonon accumulation, but from the diminished Joule heating due to lower electron mobility and current flow.
@@ IV Discussion
The lifetime of hot LO phonons determines the severity of their influence on device behavior. This lifetime arises from a complex interplay of factors, including ambient temperature, carrier density, and supplied power determining electron temperature. tsen1998time; tsen2006subpicosecond; Dyson_2009a; matulionis2009ultrafast; matulionis2009hot The lifetime is known to decrease with increasing carrier density due to plasmon–LO phonon coupling, which enhances lattice anharmonicity. dyson2008phonon The resulting coupled phonon–plasmon modes possess higher group velocities, allowing phonons to escape from the active channel region more efficiently. In bulk GaN, the lifetime of the coupled LO-phonon-plasmon mode is several 100s of femtoseconds, while at low carrier densities the LO phonon mode can live for a few picoseconds. In contrast, in AlGaN/GaN the involved wavevectors in electron-phonon interaction result in Landau damped plasma modes, making quantitative assessment of the lifetime more challenging. Ridley_Dyson_2009; dyson2011lifetime
@@
Figure 12: LO phonon lifetime reported in previous works.  tsen1998time; tsen2006subpicosecond; shi2004raman; beechem2008temperature; ramonas2003monte; PhysRevB.68.035338; matulionis2009plasmon; liberis2014hot
FKT simulations suggest that only ultrashort LO phonon lifetimes, below 40 fs, can reproduce the measured device characteristics. Longer lifetimes lead to a pronounced hot-phonon bottleneck that contradicts experimental observations. A key question is whether the measured DC characteristics could be reproduced by adjusting other device parameters in a way that would permit longer LO phonon lifetimes (e.g., $\tau_{\mathrm{LO}}\sim 100$ fs). While certain device features, such as the surface-state densities in the source and drain access regions, can be varied within physically reasonable ranges and affect the current levels to achieve a reasonable match for $\tau_{\mathrm{LO}}$ below 40 fs, the realistic variability of other model parameters is rather limited due to the fully physical basis of FKT. Device features such as the Schottky barrier height, polarization sheet charge density, or channel doping are tightly constrained by threshold voltage, contact resistance, and known epitaxial profiles; significant modification of these values leads to unphysical threshold voltages or incorrect subthreshold slopes. Among the device features that can be reasonably modified, none are capable of reproducing the measured I–V characteristics over the entire bias space when $\tau_{\mathrm{LO}}$ exceeds $\sim 40$ fs. Our prior simulations further demonstrate that increases in $\tau_{\mathrm{LO}}$ and other combinations of device features cannot simultaneously yield the near-threshold conduction and high-field current saturation. In contrast, ultrafast LO phonon decay ($\tau_{\mathrm{LO}}\lesssim 40$ fs) consistently yields agreement across all operating regions, irrespective of small variations in pertinent device features, e.g., access-region surface traps.

### v3

@@ III.3 Electron and Phonon Temperatures
The temperature variations in Fig. 10(a) also reveal certain characteristics of acoustic phonon lattice heating. In contrast to the electron temperature profile, the lattice temperature exhibits a more gradual spatial variation, increasing smoothly along the channel due to the slower energy transfer from electrons to acoustic phonons mediated by LO phonon decay. The lattice temperature’s dependence on $\tau_{\mathrm{LO}}$ also appears to be in contrast to the electron temperature. Under dynamic equilibrium conditions the power supplies inject energy into the device as $V_{\mathrm{DS}}I_{\mathrm{D}}$, which first heats the mobile electrons through Joule heating. The hot electrons, in turn, transfer their energy to the LO phonons through scattering, which subsequently decay into acoustic phonons that ultimately diffuse heat into the heat sinks. Increasing the LO phonon lifetime, $\tau_{\mathrm{LO}}$, strengthens the hot-phonon bottleneck where the long-lived LO phonons accumulate in the channel, forming a nonequilibrium population that repeatedly reabsorbs energy from the hot electrons. This elevates $T_{\mathrm{LO}}$, suppresses electron mobility, and reduces the overall current and power dissipation $V_{\mathrm{DS}}I_{\mathrm{D}}$. The resulting drop in $V_{\mathrm{DS}}I_{\mathrm{D}}$ leads to a lower rate of energy that needs to be carried away by the acoustic phonons, thereby reducing the maximum $T_{\mathrm{A}}$. This establishes a negative feedback mechanism, where we observe that increased $\tau_{\mathrm{LO}}$ leads to stronger electron heating but weaker lattice self-heating across the GaN channel. This explains the opposing trends of the peak lattice temperature and LO phonon temperature observed in Fig. 11, where increasing $\tau_{\mathrm{LO}}$ causes the peak $T_{\mathrm{LO}}$ to rise while the peak $T_{\mathrm{A}}$ decreases. Therefore, reduced lattice self-heating with longer phonon lifetime originates not directly from phonon accumulation, but from the diminished Joule heating due to lower electron mobility and current flow.
@@ IV Discussion
The lifetime of hot LO phonons determines the severity of their influence on device behavior. This lifetime arises from a complex interplay of factors, including ambient temperature, carrier density, and supplied power determining electron temperature. Tsen et al. (1998, 2006); Dyson (2009); Matulionis et al. (2010); Matulionis and Morkoç (2009) The lifetime is known to decrease with increasing carrier density due to plasmon–LO phonon coupling, which enhances lattice anharmonicity. Dyson and Ridley (2008) The resulting coupled phonon–plasmon modes possess higher group velocities, allowing phonons to escape from the active channel region more efficiently. In bulk GaN, the lifetime of the coupled LO-phonon-plasmon mode is several 100s of femtoseconds, while at low carrier densities the LO phonon mode can live for a few picoseconds. In contrast, in AlGaN/GaN the involved wavevectors in electron-phonon interaction result in Landau damped plasma modes, making quantitative assessment of the lifetime more challenging. Ridley and Dyson (2009); Dyson and Ridley (2011)
@@
Figure 12: LO phonon lifetime reported in previous works.  Tsen et al. (1998, 2006); Shi et al. (2004); Beechem and Graham (2008); Ramonas et al. (2003); Matulionis et al. (2003, 2009); Liberis et al. (2014)
FKT simulations suggest that only ultrashort LO phonon lifetimes, below 40 fs, can reproduce the measured device characteristics. Longer lifetimes lead to a pronounced hot-phonon bottleneck that contradicts experimental observations. A key question is whether the measured DC characteristics could be reproduced by adjusting other device parameters in a way that would permit longer LO phonon lifetimes (e.g., $\tau_{\mathrm{LO}}\sim 100$ fs). While certain device features, such as the surface-state densities in the source and drain access regions, can be varied within physically reasonable ranges and affect the current levels to achieve a reasonable match for $\tau_{\mathrm{LO}}$ below 40 fs, the realistic variability of other model parameters is rather limited due to the fully physical basis of FKT. Device features such as the Schottky barrier height, polarization sheet charge density, or channel doping are tightly constrained by threshold voltage, contact resistance, and known epitaxial profiles; significant modification of these values leads to unphysical threshold voltages or incorrect subthreshold slopes. Among the device features that can be reasonably modified, none are capable of reproducing the measured I-V characteristics over the entire bias space when $\tau_{\mathrm{LO}}$ exceeds $\sim 40$ fs. Our prior simulations further demonstrate that increases in $\tau_{\mathrm{LO}}$ and other combinations of device features cannot simultaneously yield the near-threshold conduction and high-field current saturation. In contrast, ultrafast LO phonon decay ($\tau_{\mathrm{LO}}\lesssim 40$ fs) consistently yields agreement across all operating regions, irrespective of small variations in pertinent device features, e.g. access-region surface traps.
@@
We note that quantum confinement effects are not included in our simulations. This choice is motivated by the nature of carrier transport in these devices, which is dominated by hot electrons in and around the pinch-off region. In this region, the electric field perpendicular to the channel–barrier interface is reduced, resulting in weaker quantum confinement. Moreover, very hot electrons, such as those present near pinch-off, can more readily access the full three-dimensional momentum space during scattering events. Under these conditions, treating channel electrons using a three-dimensional description constitutes a reasonable approximation. Using this description, FKT has successfully captured the physics underlying the experimentally observed kink effect Grupen (2019) and gate lag Miller et al. (2022) in GaN HEMTs. Moreover, most ensemble Monte Carlo–based transport simulations of GaN HEMTs rely on a similar approximation, without explicitly resolving carrier wave functions from Schrödinger’s equation. Fang et al. (2019); García-Sánchez et al. (2025)
Finally, we briefly comment on our treatment of optical phonons. We assume a $\mathbf{q}$-symmetric optical phonon distribution, with the upper bound on $|\mathbf{q}|_{\mathrm{max}}$ determined from bulk GaN simulations, as an effective description of hot-phonon accumulation. Varying $|\mathbf{q}|_{\mathrm{max}}$ within a reasonable range $(5-7)\times 10^{6}$ cm-1 results in only minor quantitative differences and does not alter the main conclusions of this work. In addition, within our solver all hot electrons are allowed to emit phonons: for a hot electron with a given initial wave vector $\mathbf{k}$ all possible final states $\mathbf{k}^{\prime}$ are considered. We find that most LO phonons are emitted by electrons in the drain access region adjacent to the pinch-off region. Owing to the high electron density in this region, the asymmetry in the electron distribution is expected to be relatively small, suggesting only a weak preference for forward versus backward optical phonon emission. Consequently, while anisotropic phonon distributions may introduce quantitative corrections, the primary conclusions of the paper remain robust.